# Stratified Expert Cloning with Adaptive Selection for User Retention in Large-Scale Recommender Systems

### v1

@@ 5.1.3. RQ1 Results
Table 1 presents the comprehensive offline experimental results, comparing the performance of SEC against various baseline methods. SEC demonstrates superior performance across most metrics, particularly in terms of Return Time, Click Rate, and Long View Rate. Focusing on the Return Time metric, which directly reflects user retention, SEC achieves a significant reduction of 5.7% compared to the best baseline model, GFN. This improvement indicates that SEC effectively prolongs user lifecycles and encourages users to return to the platform more frequently. Similarly, SEC outperforms GFN in terms of Click Rate and Long View Rate, with improvements of 3.5% and 3.9%, respectively. These results suggest that SEC not only enhances user retention but also increases user engagement, as evidenced by higher click-through rates and longer view durations.
However, it is worth noting that SEC’s performance in terms of Like Rate is lower compared to GFN, with a difference of 8.9%. This can be attributed to the primary objective of SEC, which is to optimize long-term user retention. In pursuit of this goal, SEC may prioritize recommendations that encourage users to stay on the platform and explore more content, rather than solely focusing on maximizing immediate user satisfaction, as indicated by the Like Rate.
It is important to recognize that user retention and immediate user satisfaction are distinct objectives, and there may be trade-offs between them. While a high Like Rate is generally desirable, it does not necessarily translate to long-term user retention. Users may express immediate satisfaction with a recommendation by liking it, but this does not guarantee their continued engagement with the platform over an extended period.
5.1.4. RQ2 Results
We conduct an ablation study to analyze the contribution of each component in SEC. Table 2 presents the results. Removing the multi-level behavior cloning leads to a 0.9% increase in Return Time and a over 1.0% decrease in both Click Rate and Long View Rate, indicating the importance of learning from experts at different retention levels in improving user retention and engagement. When removing the action entropy regularization, we observe a 1.8% increase in Return Time, similar with the overall goal of SEC to optimize long-term user retention. However, the removal of action entropy regularization also results in a 1.7% increase in Click Rate, accompanied by a 2.4% decrease in Long View Rate, suggesting that while the absence of action entropy regularization may lead to higher immediate user engagement, it may not necessarily translate to sustained engagement.
To further investigate the impact of the balance factor between behavior cloning and action entropy regularization, we conduct experiments with different values of the balance factor and evaluate their effect on the return time. Figure 6 presents the results, revealing an interesting pattern. When the balance factor is set to a low value, indicating a lower emphasis on action entropy regularization, the return time is relatively high, around 1.45. As the balance factor increases, the return time initially improves, reaching its optimal value of 1.411 at a moderate balance factor setting. However, when the balance factor becomes too large, prioritizing action entropy regularization over behavior cloning, the performance deteriorates significantly, with the return time increasing to 1.58. This observation highlights the importance of striking the right balance between behavior cloning and action entropy regularization in SEC.
To further investigate the impact of action entropy regularization, we visualize the learned action representations using t-SNE, as shown in Figure 5. The visualization reveals that the action embeddings learned by SEC exhibit significant diversity, with a high degree of entropy. Rather than forming distinct clusters, the actions are spread out in the embedding space, indicating that SEC learns a wide range of recommendation strategies.
5.2. Online Experiments
Figure 7. Online experimental environment, consistent with GFN (Liu et al., 2024a), where the SEC model collaborates with other strategies to optimize user retention.
5.2.1. Experimental Setup
In the online experiments, we deploy SEC on two major video platforms, Kuaishou and Kuaishou Lite, with hundreds of millions of daily active users. In the offline experiments, our strategy has full control over the ranking of items, allowing the actions to be combinations of multiple predictive factors, resulting in a continuous action space. Nevertheless, as demonstrated in Figure 7, during online experiments, we must incorporate the recommendation rankings of other scores that are beyond our control. Given the large scale of video IDs, directly using them as actions would be computationally infeasible. Therefore, we employed video clustering. Specifically, we clustered videos based on multimodal features, grouping similar videos together. The recommendation actions are then generated based on these clusters, which not only improves computational efficiency but also enhances the diversity and relevance of the recommendations. State Encoder and Action Predictor both use MLP network. We used users with high activity days as experts and classify them into 3 levels. During online recommendation, we combine cluster action prediction probability with other fine-grained ranking scores, which already include GFN (Liu et al., 2024a) as a baseline, and apply them in the ranking stages, as shown in Figure 7.
We conduct online A/B tests, optimizing for long-term user retention. In the tests, we allocate 10% of the real traffic to the baseline version and the optimized version (using SEC) and run the experiments for over two weeks.
5.2.2. RQ3 Results
The experiments focus on user retention, and the results are presented in Table 3. We use the LT7 metric to calculate Active Days, defined as the number of active users within a 7-day window divided by the total number of users in the group.
@@ 5.2.2. RQ3 Results
Table 3. Online A/B test results for User Retention. Bold values indicate statistically significant results ($p$-value $<5\%$).
Model Stage Kuaishou Kuaishou Lite
Behavior Cloning + Entropy Loss 1st Ranking +0.034% +0.037%
@@
The experimental results demonstrate that SEC achieves significant improvements in Active Days across both platforms. Initially, we implemented a model with Behavior Cloning + Entropy Loss, which showed promising results. For the 1st Ranking stage, Kuaishou showed a gain of +0.034%, while Kuaishou Lite exhibited an improvement of +0.037%. In the 2nd Ranking stage, these gains increased to +0.036% for Kuaishou and +0.042% for Kuaishou Lite.
After observing these positive results, we enhanced our approach by incorporating the Multi-level Expert stratification component of SEC, replacing the previous model in both ranking stages. This further improvement led to additional gains of +0.028% on Kuaishou and +0.043% on Kuaishou Lite.
Overall, the cumulative improvements are +0.098% for Kuaishou and +0.122% for Kuaishou Lite, additionally bringing tens of thousands of daily active users to each platform. These results validate the effectiveness of the complete SEC framework in prolonging user lifecycles and enhancing user retention.
Additionally, we examine the impact of incorporating action entropy regularization into our training process. The results indicate that this addition improves AUC scores during online training, with varying levels of improvement observed across different activity levels of experts, as shown in Figure LABEL:fig:auc_results. Notably, our AUC calculations are based on a dataset containing billions of samples. This suggests that our action entropy regularization and behavior cloning loss are isotropic, effectively enhancing overall model performance.
@@ 5.2.3. RQ4 Results
To further demonstrate the effectiveness of our approach, we conduct an additional set of experiments focusing on user interest expansion. The results are presented in Table 4.
Table 4. Online A/B test results for User Interest Expansion. Bold values indicate statistically significant results ($p$-value $<5\%$).
@@
We evaluate the impact of SEC on the number of valid interest clusters (Valid Clusters) for users. Valid interest clusters refer to the video clusters that users frequently interact with within a certain time period. This metric indirectly reflects the breadth of user interactions and the diversity of their interests on the platform. The experimental results show that SEC increases the Valid Clusters metric by 1.31% on Kuaishou and 1.14% on Kuaishou Lite.
The online experiments highlight the practical value of SEC in industrial-scale recommendation systems. By learning from the behaviors of experts at different retention levels, adaptively selecting the best expert strategies, and encouraging diversity in recommendation actions, SEC achieves notable improvements in both user retention and user interest expansion, depending on the optimized long-term value. The use of clustering to handle the vast scale of video content not only reduces computational complexity but also enhances the generalizability and robustness of the recommendation system, providing valuable insights for engineering practices in large-scale recommendations.
@@ 6. Conclusion
We propose Stratified Expert Cloning (SEC), a novel imitation learning framework for optimizing long-term user retention in large-scale recommender systems. SEC leverages multi-level behavior cloning to learn specialized policies for expert users at different retention levels, an adaptive expert selection mechanism to assign users to the most appropriate level, and action entropy regularization to promote recommendation diversity and exploration.
Extensive offline experiments demonstrate SEC’s superiority over state-of-the-art methods in improving retention metrics. Our online A/B tests on two major video platforms, Kuaishou and Kuaishou Lite, validate the practical impact of our approach with cumulative lifts of 0.098% and 0.122% in active days respectively, additionally bringing tens of thousands of daily active users to each platform. Our ablation studies confirm the effectiveness of each SEC component, with the multi-level expert stratification providing significant additional gains beyond basic behavior cloning with entropy regularization.
This work highlights the promise of imitation learning for retention-aware recommender systems in production environments. Future directions include incorporating inverse reinforcement learning, semi-supervised learning, and transfer learning. SEC lays a foundation for developing recommenders that deliver long-lasting user value at scale, with proven business impact across multiple platforms.

### v2

@@ 5.1.3. RQ1 Results
Table 1 shows SEC outperforms baselines across key metrics. For Return Time, SEC achieves a 5.7% reduction compared to GFN, indicating improved user retention. SEC also surpasses GFN in Click Rate and Long View Rate by 3.5% and 3.9% respectively, demonstrating enhanced user engagement.
However, SEC’s Like Rate is 8.9% lower than GFN’s, reflecting a trade-off between immediate satisfaction and long-term retention. This aligns with SEC’s primary objective of optimizing user retention rather than maximizing immediate satisfaction.
@@
Table 3. Online A/B test results for User Retention. Bold values indicate statistically significant results (p-value ¡ 5%). The total gains translate to over 200,000 additional daily active users (DAU) for each platform.
Metric Model Stage Kuaishou Kuaishou Lite
Active Days Behavior Cloning + Entropy Loss 1st Ranking +0.034% +0.037%
@@
5.1.4. RQ2 Results
Our ablation study (Table 2) shows removing multi-level behavior cloning increases Return Time by 0.9% and decreases Click Rate and Long View Rate by over 1.0%, demonstrating the importance of learning from experts at different retention levels.
Removing action entropy regularization increases Return Time by 1.8% and Click Rate by 1.7%, but decreases Long View Rate by 2.4%. This suggests that while entropy regularization may reduce immediate engagement, it enhances sustained engagement.
Figure 6 shows SEC’s performance with different balance factors. The return time is optimized at moderate values (around 0.01), while very high or low values lead to performance degradation. This highlights the importance of balancing behavior cloning and action entropy regularization.
5.2. Online Experiments
@@
5.2.1. Experimental Setup
We deployed SEC on Kuaishou and Kuaishou Lite platforms, each with over 200 million daily active users. Unlike offline experiments where our strategy controlled item ranking completely, in online environments, we integrate with existing ranking systems.
Due to the vast scale of video IDs, we used video clustering based on multimodal features to generate recommendation actions. We combined cluster action prediction probability with other fine-grained ranking scores in the two ranking stages to rank videos. We employed MLP networks for both State Encoder and Action Predictor, with users having high activity days classified into 3 expert levels. In our A/B tests, we allocated 10% traffic to both baseline and SEC-optimized versions for over two weeks.
5.2.2. RQ3 Results
Table 3 shows SEC significantly improved Active Days (users active within a 7-day window) across both platforms. First, we implemented Behavior Cloning + Entropy Loss, achieving +0.034% and +0.037% gains in 1st Ranking stage, and +0.036% and +0.042% in 2nd Ranking stage for Kuaishou and Kuaishou Lite respectively.
Adding Multi-level Expert stratification yielded additional gains of +0.028% and +0.043%. The cumulative improvements (+0.098% for Kuaishou, +0.122% for Kuaishou Lite) translate into over 200,000 additional daily active users on each platform, underscoring SEC’s real-world impact at scale.
Additionally, we examine the impact of incorporating action entropy regularization into our training process. The results indicate that this addition improves AUC scores during online training, with varying levels of improvement observed across different activity levels of experts, as shown in Figure 7. Notably, our AUC calculations are based on a dataset containing billions of samples. This suggests that our action entropy regularization and behavior cloning loss are isotropic, effectively enhancing overall model performance.
@@ 5.2.3. RQ4 Results
We also evaluated SEC’s impact on user interest expansion (Table 4). Valid interest clusters (video clusters users frequently interact with) increased by 1.31% on Kuaishou and 1.14% on Kuaishou Lite, demonstrating SEC’s ability to broaden user interests.
These results highlight SEC’s practical value in industrial-scale recommendation systems, showing improvements in both user retention and interest diversity through learning from experts at different retention levels, adaptive selection, and encouraging recommendation diversity.
Table 4. Online A/B test results for User Interest Expansion. Bold values indicate statistically significant results ($p$-value $5\%$).
@@
6. Limitations
While the proposed SEC framework demonstrates strong performance, its applicability is constrained by several limitations. First, its effectiveness is fundamentally contingent on a sufficiently large and diverse set of high-retention ”expert” users, which may not be available on nascent platforms or in niche domains. Furthermore, the adaptive expert selection mechanism relies on high-quality user state representations; these can be sparse or unreliable for new users, posing a significant challenge in cold-start scenarios. Lastly, although validated on short-video platforms, the generalizability of SEC to other domains such as e-commerce or news recommendation warrants further investigation, as the definitions of expert behavior and retention metrics can vary significantly across contexts.
@@ 7. Conclusion
In this work, we introduced Stratified Expert Cloning (SEC), a novel imitation learning framework designed to optimize long-term user retention in large-scale recommender systems. SEC integrates three key innovations: multi-level expert stratification to capture diverse retention behaviors, adaptive expert selection to match users with appropriate policies, and action entropy regularization to enhance recommendation diversity. Our comprehensive evaluation, featuring extensive offline experiments and large-scale A/B tests, confirms SEC’s superior performance over state-of-the-art methods, with ablation studies validating the contribution of each component. By circumventing the exploration risks and delayed reward challenges inherent in reinforcement learning, SEC offers a practical and effective approach to building recommender systems that deliver long-lasting user value at scale.
Appendix A GenAI Usage Disclosure
During the preparation of this work, GenAI tools were used solely to improve the spelling and grammar of the author-written text. The authors are fully accountable for the content.